5,000 10,000 or 15,000 for drop box 2EQuestion Help Problem 2 Eight Flags operates several amusement parks in the Midwest. The company stocks machine oil to servi
Answer #1

Since Lead Time is in week, we will have to find the weekly demand as well to calculate the demand during the lead time.

Weekly Demand d = 30000/50 = 600

We will find the total cost for each alternative -

Formula for Total Annual Cost is as mentioned below -

Total annual cost = Material costs + Freight cost + Inventory cost + administrative costs

TC = p*D + Freight costs + (Q/2 + dL)H + Administrative costs

First of all, we will find the total cost at quantity 5000, 10000, 15000 for Supplier Sharps

TC (Q = 5000) = 30000*3.90 + 4800 + ((5000/2)+(600*3))*0.78 + 4500 = 129654

TC (Q = 10000) = 30000*3.90 + 3400 + ((10000/2)+(600*3))*0.78 + 4500 = 130204

TC (Q = 15000) = 30000*3.90 + 2900 + ((15000/2)+(600*3))*0.78 + 4500 = 131654

Total cost at quantity 5000, 10000, 15000 for Supplier Winkler

TC (Q = 5000) = 30000*3.80 + 5200 + ((5000/2)+(600*4))*0.76 + 4000 = 126924

TC (Q = 10000) = 30000*3.80 + 3700 + ((10000/2)+(600*4))*0.76 + 4000 = 127324

TC (Q = 15000) = 30000*3.80 + 3400 + ((15000/2)+(600*4))*0.76 + 4000 = 128924

From Above, Winkler with order quantity 5000 provide the lowest cost. Cost will be 126924.